OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (zale101):

the opposite of addition is subtraction so 2x+2 is 2x-2 the opposite of multiplication is division so divide 2 instead of multiplying it by x-2 x-2/2 Answer: f-1(x)= x-2/2 no replace x by 4 and solve :)

OpenStudy (zale101):

now*

OpenStudy (anonymous):

im confused

OpenStudy (zale101):

what part?

OpenStudy (anonymous):

all of it lol whats the answer?

OpenStudy (zale101):

i'll write it algebraically for you your question is f(x) = 2x + 2 change f(x) to y y=2x+2 solve for x y - 2 = 2x \[\frac{ y-2 }{2 }=x\] since i already what equals to x, then i'll have to switch x and y, making it equal to y instead of x \[\frac{ x-2 }{2 }=y\] now, change y to the inverse function \[f^{-1}(x)\] \[f^{-1}(x)=\frac{ x-2 }{2 }\] now, you're question says "when x = 4" after finding the inverse of the orginial equation, rreplace x with 4 \[f^{-1}(4)=\frac{ (4)-2 }{2 }\]

OpenStudy (anonymous):

ohhhh thank you that helped alot

OpenStudy (zale101):

No Problem :) Glad it did!